OpenStudy (anonymous):

I WILL GIVE A MEDAL! The tenth, eleventh, and twelfth terms of a sequence are shown in the table below: Term number 10 11 12 Term 28 32 36 Which of the following shows the first five terms of the sequence? -8, -4, 0, 4, 8 8, 4, 1, -4, -8 -4, 0, 4, 8, 12 4, 0, -4, -8, -12

OpenStudy (anonymous):

@Nnesha I NEED YOU PLEASE

OpenStudy (anonymous):

i think 1st one (10+4)*2=28 (11+5)*2=32 (12+6)*2=36 (9+3)*2= (8+2)*2= (7+1)*2= (6+0)*2= (5-1)*2=8

OpenStudy (anonymous):

ok, let me just see what nnesha says

OpenStudy (anonymous):

thank you btw

OpenStudy (anonymous):

(4-2)*2=4 (3-3)*2=0 (2-4)*2=-4 (1-5)*2=-8

Nnesha (nnesha):

okay so try to figure out how you ther number are increasing or decreasing first term is 28 and then second is 32 what you have to do with 28 to get 32 ??

OpenStudy (anonymous):

i dont know

Nnesha (nnesha):

do you have to add or subtract something into 28 to get 32 ??

OpenStudy (anonymous):

ill just stick with the first one

OpenStudy (anonymous):

do you think you can help me with one more question?

Nnesha (nnesha):

if you have 28 dollar what should u do to get 32 gain some dollar or lose some dollar ??

OpenStudy (anonymous):

gain

Nnesha (nnesha):

yes so 28 + what = 32 ??

Nnesha (nnesha):

in other way 32 -28 = ??

OpenStudy (anonymous):

4

Nnesha (nnesha):

yes so 28 +4 = 32 32+4 =36 so you are adding 4 to get next term

Nnesha (nnesha):

i'll try my best for next question :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Choose the system of equations which matches the following graph. a line includes points 0 comma 0 and negative 1 comma 2. A line includes points 2 comma 3 and negative 1 comma negative 2 5x − y = 7 2x − y = 0 5x − y = 7 2x + y = 0 5x + y = −7 2x + y = 0 5x + y = −7 2x − y = 0

Nnesha (nnesha):

\[\huge\ (0,0)(-1,2)\]first two points where are graph ??

Nnesha (nnesha):

where both lines intersect that point is your solution so look at the graph and see where bot lines intersects

OpenStudy (anonymous):

hope this helps

Nnesha (nnesha):

hmm what do you mean this help :D look at graph and see at what point both lines intersect